OpenStudy (anonymous):

eek please someone help me with this elimination problem i have no clue at all 5x-4y=14 3x+3y=3

OpenStudy (anonymous):

solve by addition (elimination)

OpenStudy (anonymous):

solve for y?

OpenStudy (anonymous):

and x

OpenStudy (anonymous):

ok 5(0)-4y=14

OpenStudy (anonymous):

5x-4(0)=14

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

and why is there two =s on the second one

OpenStudy (anonymous):

is that not what your doing? I'm sorry it looked like what i learned

OpenStudy (anonymous):

\[5x-4y=14\]\[ 3x+3y=3\]multiply the first equation all the way across by 3, the second by 4

OpenStudy (anonymous):

\[3(5x-4y)=3\times 14\]\[15x-12y=42\] for the first one

OpenStudy (nubeer):

5x-4y=14 multiply it with 3 3x+3y=3 multiply with 5 15x-12y = 42 15x+15y= 15 - - - 0x -3y = 27 y = -9 just put this value of y in any equation you will get the value of x

OpenStudy (anonymous):

\[4(3x+3y)=4\times 3\]\[12x+12y=12\] for the second one

OpenStudy (anonymous):

put them on top of each other and add in a column, get \[15x-12y=42\]\[12x+12y=12\] add in a column get \[27x=54\] as the \(y\)'s add up to zero

OpenStudy (anonymous):

therefore \(x=\frac{54}{27}=2\) now that you have \(x=2\) you can replace \(x\) by \(2\) in either equation and solve for \(y\)

OpenStudy (anonymous):

btw the second equation can be rewritten, by dividing both sides by 3, as \[x+y=1\] then you could write \(y=1-x\) and replace the \(y\) in the first equation by \(1-x\) to solve \[5x-4(1-x)=14\] for \(x\) you will get the same answer